OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2

OpenStudy (anonymous):

infinitely many solutions for both x and y

OpenStudy (ash2326):

4x+8y=8 x+2y=2 dividing first equation by 2 x+2y=2 both equation are same, so there are infinite solutions to both x and y